OpenStudy (anonymous):

how do you solve x^2-14x+49=9 by using the square root property?

myininaya (myininaya):

\[\text{ recall } x^2-14x+49=(x-7)^2\]

myininaya (myininaya):

use that and have fun

myininaya (myininaya):

look at the previous one i did for you and ask any questions but you should try this one on your own and i will totally check if you like

OpenStudy (anonymous):

Notice how x^2, 9 and 49 are perfect squares? This will come in handy: First, factor the left hand side by looking for two numbers that add up to -14 and multiply to 49 (these two numbers are -7 and -7) so you get \[(x-7)^2=9 \] square root both sides and solve for x: \[x-7=3\]\[x=10\]

myininaya (myininaya):

plus or minus 3 right junkie?

OpenStudy (anonymous):

Yeah, just noticed that there should be two solutions

myininaya (myininaya):

\[x-7=\pm 3\]

OpenStudy (anonymous):

why is it -14 thouqh ?

OpenStudy (anonymous):

oh nevermind iget why . THankyou !

OpenStudy (anonymous):

well when you factor a quadratic function you generally have the form : \[x^2+ax+b\]and you look for two numbers (c & d) that satisfy the following two equations:\[c+d=a\]\[c*d=b\]c and d then get plugged into this guy: \[(x+c)(x+d)\]

OpenStudy (anonymous):

Ohhhhh okay ! :D